A car gets 20 kms per gallon of gas when it travels at 50 kms per hour. The car gets 12% fewer kms to the gallon at 60 kms per hour. How far can the car travel at 60 kms per hour on 11 gallons of gas?
Answer by ikleyn(52834)

.
A car gets 20 kms per gallon of gas when it travels at 50 kms per hour.
The car gets 12% fewer kms to the gallon at 60 kms per hour.
How far can the car travel at 60 kms per hour on 11 gallons of gas?
~~~~~~~~~~~~~~~~~~~~

At the speed of 60 km/h, the car travels 20*(1-0.12) = 20*0.88 = 17.6 kilometers per gallons.


Using 11 gallons of gas, the car can travel 11*17.6 kilometers = 193.6 kilometers.  ANSWER

Solved.